OpenStudy (anonymous):

MEDAL AND FAN what is the simplified form of the following expression? (5/4)^3

Nnesha (nnesha):

hint:\[\huge\rm (\frac{ a }{ b })^m =\frac{ a^m }{ b^m }\]

OpenStudy (anonymous):

\[(\frac{ 5 }{ 4 })^{3}=\frac{ 125 }{ 64 }\]